Rice a costs PHP36/kilo and rice B costs PHP32/kilo.How many kilos of each kind must be mixed to produce a 30 kilos mixture which costs PHP33.60/kilo?
Answer by jorel1380(3719)

Let n be the amount of rice A, in kilos. Then rice B would be 30-n. So:
36n+32(30-n)=33.6(30)
4n+960=1008
4n=48
n=12
You need 12 kilos of rice A; and 18 kilos of rice B for your desired mixture. ☺☺☺☺